OpenStudy (anonymous):

which conditonal and converse are both true? if x =2 then x^2=4 if x-1 then 2x=2 if x=3 then x^2=6 If x^2 =4 then x=2

OpenStudy (loser66):

the second one should be x =1 , right?

OpenStudy (blurbendy):

The last one works, because if x^2 = 4, then logically, x = 2 the converse: If x = 2, then x^2 = 4 is true as well

OpenStudy (loser66):

I don't think so,

OpenStudy (anonymous):

yes sorry

OpenStudy (loser66):

the second one, because "if " is condition and converse is backward that means the conclusion turns to " if". so if x =1 then 2x =2 (true) converse: if 2x =2 then x =1 (true) therefore the second one is correct one

OpenStudy (blurbendy):

i think Loser66 is right because x^2 = 4 would have two answer +2 and -2. good catch

OpenStudy (blurbendy):

answers*